Suppose that King Kanuta from Problem 11 demands that each of his subjects give him 4 coconuts for every coconut that they consume. The king puts all of the coconuts that he collects in a large pile and burns them. The supply of coconuts is given by S(ps) = 100ps, where ps is the price received by suppliers. The demand for coconuts by the king’s subjects is given by D(pd) = 10,400 - 100pd, where pd is the price paid by consumers. In equilibrium, the price received by suppliers will be

a. $20. b. $52. c. $30. d. $260. e. None of the above.
Answer #1

Equilibrium is when demand is equal to supply

100p = 10400 -100p

Which implies 200p = 10400

which implies p= 10400/200 = $52

Therefore (b) $52 is the answer to the question